Name:

SS#:

Test 2

Instructions: Answer all six problems, two on each answer sheet. Put the number of the problems you are doing at the top of the answer sheet.

All problems are worth the same amount.

1. 40% of all Americans prefer to fix their eggs scrambled. 8 Americans are selected at random.

(a) What is the probability that at most two prefer their eggs scrambled?

(b) What is the expected nmber of people who want their eggs scrambled?

2. Assuming the graphs below correspond to the standard normal distribution Z, find the values z.

3. The distance students can throw their statistics textbook is normally distributed with mean 30’, and standard deviation 5’.

(a) What is the probability that a random person throws their book a distance between 25’ and 33.5’?

(b) What is the probability that they throw it over 42’?